In a test of the effectiveness of garlic for lowering cholesterol 51 subjects were treated with raw garlic. Cholesterol levels were measured before and after the treatments. The changes (before and after) in the levels of LDL cholesterol (in mg/dL) had a mean of 0.4 and a sample standard deviation of 21. Construct a 98% confidence interval estimate of the mean net change.
Answer #1

8012 n=51 Standard mean co.4 deviation = 21 98 % confidence interval: = (a + tcxsin) = 60.4+ 2.403 *) = 10.4 + 700662) = 6-61